6.1-4 A signal g(t) = sinca (571) is sampled (using uniformly spaced impulses) at a rate of (1) 5 Hz, (ii) 10 Hz; (ii) 20 Hz. solve it using matlab.. show the code
Answer #1

If you have any doubts, please comment. I will definitely clarify it. Please give positive rating for my efforts. Thanks.

CODE for 5 Hz

%% for Fs = 5hz (sampling frequency) close all;clear all;clc; Fs = 5; %sampling freq N = 51; t = 0:1/Fs:(N-1)/Fs; x = sin(5*pi.*t).^2; Nf = 1024; X = fftshift(fft(x,Nf)); f = (-Fs/2:Fs/Nf:Fs/2-Fs/Nf); figure(1) subplot(311) stem(t,x) xlabel("time") ylabel("amplitude") subplot(312) plot(f,abs(X).^2/Nf) xlabel("frequency") ylabel("magnitude") subplot(313) plot(f,angle(X)*180/pi) xlabel("angle") ylabel("magnitude")

IMAGE OF CODE

| | data_extract.m x BCI-IV.m x sampling. m x sam 8% for Fs = 5hz (sampling frequency) close all;clear all;clc; Fs = 5; $samp

IMAGE OF OUTPUT

Figure 1 - O X Help File Edit View Insert Tools Desktop Window dua @ IE E amplitude anekulelohet 09 Yleliolella on 7 8 9 10 t

CODE for 10 Hz

%% for Fs = 10hz (sampling frequency) close all;clear all;clc; Fs = 10; %sampling freq N = 51; t = 0:1/Fs:(N-1)/Fs; x = sin(5*pi.*t).^2; Nf = 1024; X = fftshift(fft(x,Nf)); f = (-Fs/2:Fs/Nf:Fs/2-Fs/Nf); figure(1) subplot(311) stem(t,x) xlabel("time") ylabel("amplitude") subplot(312) plot(f,abs(X).^2/Nf) xlabel("frequency") ylabel("magnitude") subplot(313) plot(f,angle(X)*180/pi) xlabel("angle") ylabel("magnitude")

IMAGE OF CODE

| data_extract.m x BCI-IV. m x sampling1.m x sampli for Fs = 10hz (sampling frequency) close all;clear all;clc; Fs = 10; $sam

IMAGE OF OUTPUT

- 0 X Figure 1 File Edit View 3 3 Insert Window Help Tools Desktop E A E 1994-9-99999999994-9 -14 TIITTI amplitude رلهلهلهلهل

CODE for 20 Hz

%% for Fs = 20hz (sampling frequency) close all;clear all;clc; Fs = 20; %sampling freq N = 51; %number of points t = 0:1/Fs:(N-1)/Fs; x = sin(5*pi.*t).^2; Nf = 1024; X = fftshift(fft(x,Nf)); f = (-Fs/2:Fs/Nf:Fs/2-Fs/Nf); figure(1) subplot(311) stem(t,x) xlabel("time") ylabel("amplitude") subplot(312) plot(f,abs(X).^2/Nf) xlabel("frequency") ylabel("magnitude") subplot(313) plot(f,angle(X)*180/pi) xlabel("angle") ylabel("magnitude")

IMAGE OF CODE

| | | data_extract.m x BCI-IV.m x sampling1.m x samp %% for Fs = 20hz (sampling frequency) close all;clear all;clc; Fs = 20;

IMAGE OF OUTPUT

- O X Figure 1 File Edit View Insert Tools Desktop Ha @ DE E Window Help amplitude time magnitude -10 8 6 4 2 2 4 6 8 10 0 fr